OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!! Solve the inequality: 2t + 1 > 13 or –18 > 7t + 3

OpenStudy (anonymous):

@Vocaloid @pooja195 @Peaches15

OpenStudy (anonymous):

@absolutemaster

OpenStudy (anonymous):

@hype.child

OpenStudy (anonymous):

give me a second

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@ValentinaT

pooja195 (pooja195):

2t + 1 = 13 –18 = 7t + 3 Now solve both for t

pooja195 (pooja195):

2t + 1 = 13 subtract 1 from both sides 13-1=?

OpenStudy (anonymous):

ok hang on a sec

OpenStudy (anonymous):

for the first one t=6 for the second one t=-3

OpenStudy (anonymous):

is that correct?

pooja195 (pooja195):

yes then simply remove the equal signs and you get as a final answer: t>6 or t<−3

OpenStudy (anonymous):

thank you very much @pooja195

pooja195 (pooja195):

:)

OpenStudy (anonymous):

@pooja195 i have another question

OpenStudy (anonymous):

nvm